OpenStudy (anonymous):

2. Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form.(Points : 1) y = 2x y = -2x y = 1/2 x y = -1/2 x 3. Find the direct variation equation which passes through (0, 0) and (4, 1).(Points : 1) y = 4x y = -4x y = 1/4 x y = -1/4 x

OpenStudy (blurbendy):

(-2 - 0) = -2 (1 - 0) = 1 y = -2x 2) (1 - 0) = 1 (4 - 0) = 4 y = 1/4x